A tetraphenyl phosphonium chloride (TPPCl) powder (FW=342.39) is 89.0 percent pure. How many grams are needed to prepare 0.40 L of a 28.0 mM solution?
Answer #1

volume , V = 0.4 L


use:
number of mol of TPPCl,
n = Molarity * Volume
= 2.8*10^-2*0.4
= 1.12*10^-2 mol
use:
mass of TPPCl,
m = number of mol * molar mass
= 1.12*10^-2*3.424*10^2
= 3.835 g

This is mass required if purity is 100 %

use:
% purity = pure mass * 100 / impure mass
89.0 = 3.835 * 100 / impure mass
impure mass = 4.31 g
Answer: 4.31 g